1a. In the long-run, if existing firms are earning economic losses, this implies:

  • ATC is greater than price.

  • AFC is greater than price.

  • AFC is equal to price.

  • ATC is less than price.

1b.

A firm has the following price and cost information

Price $10.00
Output 20
Fixed Costs $200.00
Variable Costs $1,200.00

Instructions: Enter all prices as 2 decimal places.

Question 1:

What is the average fixed cost? $

Question 2:

What is the average variable cost? $

Question 3:

What is the average total cost? $
Answer #1

Ans) the correct option is a) ATC is greater than price.

Ans) 1) average fixed cost = 200/20 = $ 10

2) Average variable cost = 1200/20 = $ 60

3) average total cost = 10 + 60 = $ 70